OpenStudy (anonymous):

(2y+5)(y+2)=-5(2y+5)

OpenStudy (anonymous):

Divide each side by (2y+5) and you end up with: y+2 = -5 y= -7

OpenStudy (anonymous):

(3p)\[(3p )^{2}p ^{3}/p ^{-1}p ^{-5} \times(8p ^{-4})^{-3}\]

OpenStudy (anonymous):

needs to be simplified

OpenStudy (lalaly):

2+y=-5 y=-7

OpenStudy (anonymous):

the best way is to divide both sides by (2y+5) y+2=-5 y=-7

OpenStudy (anonymous):

Are the p^-1 and p^-5 both under the first set of variables?

OpenStudy (anonymous):

multiplying it all out and collecting like terms and solving for y by getting the zeros works. the roots are -5/2 and -7

OpenStudy (anonymous):

He has a better answer. I was simplifying. I would submit the last answer.